# A Survey on Data Selection for Language Models

### v2

@@ 3.1 Language Filtering
When curating data for language model pretraining, a crucial first step is to consider the languages the model will operate in and to filter out data that doesn’t belong to those languages. This applies not only to natural languages, but coding languages as well, however, the methods to determine each will differ. For multilingual language models, in addition to filtering out undesirable languages, it is also important to track metrics on the quantity of data coming from each language. Similarly, for models with code capabilities, it is important to track the quantity of data from each coding language.
@@
When filtering for language, it is crucial that the utility functions be fast to compute as the utility calculation will be performed on huge quantities of data. Thus, many methods that aim to filter for specific natural languages utilize fast-to-compute utility functions from classifiers based on character n-grams (Conneau & Lample, 2019; Wenzek et al., 2020; Raffel et al., 2020; Xue et al., 2021; Laurençon et al., 2022) including langdetect [^2] https://pypi.org/project/langdetect/, cld3333https://github.com/google/cld3, and fastText (Joulin et al., 2016; Grave et al., 2018). Some methods aim to remove all non-English data, while others may include over 100 languages and require different utility function strategies.
@@
When developing the English-only C4Raffel et al. (2020) use a naive Bayes classifier with character n-grams (langdetect) to filter out any pages not classified as English with a probability of 0.99. They find the use of a very high threshold sensitivity to be appropriate as the filtered dataset still contains a significant quantity of data ( 750GB of data). More recently, when developing DolmaSoldaini et al. (2024) utilize fastText’s (Grave et al., 2018) language identification model and filter out any documents classified as English with a probability less than 0.5, finding that this removed 61.7% of web pages. However, when filtering books (from Project Gutenberg [^4] https://www.gutenberg.org/) they first split each book by paragraph and compute the language score for each paragraph, only removing books that have an average probability of English (assigned by fastText) of below 0.5. Penedo et al. (2023) also use fastText when filtering RefinedWeb, but differently to Dolma, they use a higher threshold (0.65) and filter out higher quantities of web pages.
The lower thresholds used by the more recent methods is likely due to a combination of factors. First, when developing C4, 750Gb of data was considered a very large dataset, but more recent datasets can be well into the terabytes worth of data, and using a lower threshold will allow more data to pass through the filter. Additionally, it is difficult to measure the differences in accuracy across different language identification models (e.g. langdetect vs. cld3. vs. fastText) due to the differences in supported languages and strengths on text of varying lengths, but some experiments show that fastText has improved accuracy over other language identification models, [^5] https://modelpredict.com/language-identification-survey as well as reduced latency. In general, the threshold used for keeping or filtering out a data point can be in part determined by the desired dataset size, where a lower threshold will retain higher quantities of data, but possibly containing non-English text.
@@
The utility functions that have been used to detect code languages are very rudimentary. For example, Chowdhery et al. (2023) filter for data from 24 programming languages simply by searching for documents that match a set of approved filename extension (e.g. “.py” for python). This strategy works well when filtering over data from known code data that contains file names, such as snapshots of Github repositories (Chen et al., 2021; Li et al., 2022a). However, no filtering methods that we know of find code within natural language documents, which is entirely possible within some domains such as in Stack Exchange [^6] https://archive.org/details/stackexchange and may contribute meaningful quantities of interleaved natural language and code.
@@
There are trade-offs occurring in the filter sensitivity for language detection, where a lower threshold allows greater quantities of data to pass through the filter, at the risk of sacrificing some quality. On the one hand, using a strict filter for English data may reduce the quantities of non-English data, but on the other hand, using a lower threshold can mitigate some of the inherent biases of the language detector against dialects spoken by minority groups (e.g. African-American Vernacular English (Blodgett et al., 2016)). Determining the exact threshold to use for language classifiers is highly dependent on the use case. While fastText is the current standard for language detection due to its combination of speed and accuracy, it is not perfect. More accurate systems built in on recurrent- (Toftrup et al., 2021) and transformer-based [^7] https://huggingface.co/papluca/xlm-roberta-base-language-detection architectures have been developed, but run with higher computational costs.
Language filtering is an important first step of data selection where documents are selected to include only the desired languages. Classifier-based methods are the norm when creating both english-only and multilingual datasets. For lower-resourced languages, URL-based methods are also useful for classifying the language used in multilingual datasets and can sometimes can be more reliable than classifiers. Filtering for code languages is very simple, most works simply using the file extension. [Uncaptioned image] One interesting direction for future research is developing methods that find code data within documents of natural language. To the best of our knowledge, no such methods currently exist, but a code-language identifier could be used in domains with interleaved code and natural language to find meaningful quantities of new code data. Language Filtering Summary & Ideas
@@ 3.2 Heuristic Approaches
When pretraining a language model, raw datasets are generally composed of massive quantities of text from a wide variety of sources that are filtered through a set of simple heuristics. Major models are often trained on web scrapes such as CommonCrawl and GitHub, though transparency into their precise compositions are on the decline (Bommasani et al., 2023). It is widely known however that pretraining data contains significant quantities of boiler-plate text, error messages, and offensive text (Raffel et al., 2020; Touvron et al., 2023a). For example, Gao et al. (2020) find that, when creating the Pile, the most common 13-grams were character repetitions such as a string of dashes (“– –”) with 11 million instances. Removing such undesirable text is very important, but must be done efficiently due to the size of the corpora involved. In this case, a common approach to filtering data involves simple and efficient-to-compute heuristics.
The goal with heuristic approaches is to constrain the training distribution along some dimension (e.g. sentence length, repetitiveness), with the assumption that the evaluation distribution will exhibit similar characteristics. The number of heuristics that have been used in past works is extensive, but generally fall into one of the following categories of heuristics: item count, repetition count, existence, ratio, or statistics. In this section we discuss a representative sample of them, provide Table 2 as an overview of heuristic methods, and include a comprehensive listing in Appendix A.
@@
For heuristic-based text filtering, the utility functions should be very fast to compute and generally rely on one or more qualities of the raw text included in each data point of the dataset. However, the exact text qualities that lead to best performance depend on the desired use case for the model, and are yet unknown. Thus, a wide variety of distinct utility functions have been proposed, each making assumptions about the expected testing distribution, and filtering out data that does not fit those assumptions. For example, Raffel et al. (2020) discard any web page with fewer than five sentences, Rae et al. (2021) discard documents with fewer than 50 or greater than 100,000 words, and Xue et al. (2021) discard documents with fewer than three lines of text with 200 or more characters. Another common heuristic is to remove documents with specific black-listed words or phrases. Raffel et al. (2020) remove documents with any word on the "List of Dirty, Naughty, Obscene, or Otherwise Bad Words". [^8] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words However, Penedo et al. (2023) suggest that the NSFW word block lists as used by Raffel et al. (2020) can often lead to false positives, resulting in the over-filtering of legal and medical content.
@@
Another dimension to consider when designing a filter is the domain of interest. For example, when collecting code data for the ROOTS dataset, Laurençon et al. (2022) filter source files between 100 and 200,000 characters, with between 15-65% alphabetic characters (a-zA-Z), and line lengths between 20-1000 characters. If these filters were applied on natural language data, significant quantities of desirable data would be removed (e.g. many books have greater than 65% alphabetic characters). Additionally, Laurençon et al. (2022) filter out files which have line lengths that are highly uniform (line length standard deviation 3 characters or less). While filtering out data with uniformly lengthed lines may improve data quality for code, it would likely filter out significant quantities of poems and other rhythmic text due to the similarity in line length, and would likely remove data which has been collected from pdf files, which could be desirable when training a model for a non-code domain.
@@
Many heuristic utility functions can be used as part of a deterministic or stochastic selection mechanism. In a deterministic selection mechanism, any data points which are above/below the utility threshold will be removed. The vast majority of existing heuristic filters use a deterministic selection mechanism. However, this may unnecessarily remove data that resembles the evaluation distribution (Brown et al., 2020). In particular, the target dataset distribution may be long-tailed in the characteristic of interest, and deterministically removing all data above/below a threshold can harm the modelling of that characteristic.
@@
One disadvantage to heuristic selection approaches is that they do not judge the quality or content of a document, rather, they rely entirely on surface level statistics and counts, lacking the finesse of other data selection methods. This can lead to desirable data being removed, but this is often accepted as a trade-off as long as the quantity of data after filtering is still within the desirable range.
@@
Heuristics are a very efficient method for removing large quantities of data, but lack finesse. Most common heuristic utility functions fall into 5 categories: item count, repetition count, existence, ratio, and statistics. Heuristics should be designed specifically for the domain of interest (e.g. heuristics for code will not work for natural language). Heuristics can also be used to improve data point characteristics (e.g. removing HTML from a web document, or specific lines within a document). [Uncaptioned image] Because heuristics lack finesse, it may be worth exploring the use of stochastic selection mechanisms to allow higher quantities of data through the filter (with the hope that a filter farther down the pipeline will remove data that is actually undesirable). [Uncaptioned image] Validating the quality of a heuristic filter is a very slow process, requiring either manual inspection, or training and evaluating a model. Developing methods that can directly measure the data itself could significantly improve iteration time. Heuristic Approaches Summary & Ideas
@@ 3.3 Data Quality
Training on the highest quality data can lead to stronger performance (Du et al., 2022). However, “high-quality” is not a well defined term with respect to language model pretraining data. Additionally, there is no one-size-fits-all characterization as, depending on the use case, “high-quality” data can vary drastically (Longpre et al., 2023c). Additionally, implementations of toxicity filters and quality filters can further marginalize communities whose text is not commonly considered “high quality” (Xu et al., 2021; Longpre et al., 2023c). In this work, we narrow the use of the phrase “high-quality” to a common use of the phrase referring to: data that is known to be written by humans, and has likely gone through an editing process (Grave et al., 2018; Gao et al., 2020; Brown et al., 2020; Chowdhery et al., 2023). Some data domains that fall under the “high-quality” category are Wikipedia, books, patents, and peer-reviewed journal articles.
@@
One commonly used method when filtering for quality is classifier-based quality filtering, where the goal is to identify data points that are likely from the same (or similar) distribution as a known “high-quality” corpus of data points (reference corpus) (Longpre et al., 2023c). For example, Brown et al. (2020) filters a version of Common Crawl [^9] https://commoncrawl.org/ based on similarity to “high-quality” reference corpora (including WebText (Radford et al., 2019), Books1, Books2, and English wikipedia). They train a classifier using the “high-quality” reference corpora as the positive class and unfiltered Common Crawl documents as the negative class. This classifier then defines the utility metric for selecting data that follows the desired distribution. The classifier used to assign quality scores to each document should be fairly lightweight due to the scale of pretraining datasets. For example, Du et al. (2022) use a feature hash based linear classifier, while Brown et al. (2020) and Gao et al. (2020) use a fastText model (Joulin et al., 2016) with an n-gram size of 2. Similarly, Xie et al. (2023b) computes the utility as an importance weight between two hashed n-gram generative models. While modern language models may be able to better model the reference corpus distribution, they would also significantly increase the amount of compute required over hashing and n-gram based methods.
@@
One potential negative impact of classifier-based quality filtering is that the reference corpora are unlikely to contain examples of all data that are considered high quality. Thus, it is desirable to allow some stochasticity in the selection mechanism. For example, Brown et al. (2020) use their classifier to score Common Crawl documents and keep each document only if they satisfy
@@
where $\alpha=0.9$ was determined by matching the distribution of scores from the WebText corpus. By selecting documents that don’t match the “high quality” distribution with some stochasticity, they ensure that documents in the final corpus are mostly high-scoring, but still include some documents that are out of distribution with respect to the reference corpora.
@@
An important consideration when performing quality filtering is that certain reference corpora may be biased towards or away from certain demographics, dialects, and socialects (Rae et al., 2021). Additionally, Gururangan et al. (2022) find that a quality filter trained on Wikipedia, books, and newswire (a replication of the quality filter from Brown et al. (2020)), evaluated on high school newspaper articles, prefers articles from high schools in wealthier, higher-educated, and urban areas. They demonstrate that even though quality filtering is often assumed to be fairly neutral, it carries with it an implied judgment of preferred values. Thus, quality filtering should be done with care to ensure that biases are avoided and the distribution of data sufficiently covers the desired dialects and demographics. Filtering text for quality, while avoiding the introduction of biases is an important direction for future research.
@@
Data quality methods aim to select data points that are similar to data which is of known “high-quality”. Classifier- and perplexity-based quality filtering are the existing approaches, with classifier-based filtering being the more popular method. For both classifier- and perplexity-based methods, it is important to be careful when training the classifier because this can introduce biases (e.g. against lower-income, less-educated populations). [Uncaptioned image] Exactly which situations quality filtering is beneficial is still undecided, as recent works have trained performant models without using quality filters at all. This is an area that requires further study. Data Quality Summary & Ideas

### v3

@@ 3.1 Language Filtering
When curating data for language model pretraining, a crucial first step is to consider the languages the model will operate in and to filter out data that doesn’t belong to those languages. This applies not only to natural languages, but coding languages as well, however, the methods to determine each differ in practice. For multilingual language models, in addition to filtering out undesired languages, it is also important to track metrics on the quantity of data coming from each language. Similarly, for models with code capabilities, it is important to track the quantity of data from each coding language.
@@
When filtering for language, it is crucial that the utility functions be fast to compute as the utility calculation is often performed on all available raw text data (potentially a huge quantity). Thus, many methods that aim to filter for specific natural languages utilize fast-to-compute utility functions from classifiers based on character n-grams (Conneau & Lample, 2019; Wenzek et al., 2020; Raffel et al., 2020; Xue et al., 2021; Laurençon et al., 2022) including langdetect [^3] https://pypi.org/project/langdetect/, cld3444https://github.com/google/cld3, and fastText (Joulin et al., 2016; Grave et al., 2018). Some methods aim to remove all non-English data, while others may include over 100 languages and require different utility function strategies.
Figure 4: An overview of the data filtering pipeline for pretraining. Each filtering component is described in Section 3, and depicts common filters used for preprocessing text data. Note that different works employ different filters, at different stages, and do not necessarily adhere to the order conveyed here. This figure was adopted and modified from Soldaini et al. (2024).
@@
When developing the English-only C4Raffel et al. (2020) use a naive Bayes classifier with character n-grams (langdetect) to filter out any pages not classified as English with a probability of 0.99. They find the use of a very high threshold sensitivity to be appropriate as the filtered dataset still contains a significant quantity of data (~750GB of data). More recently, when developing DolmaSoldaini et al. (2024) utilize fastText’s (Grave et al., 2018) language identification model and filter out any documents classified as English with a probability less than 0.5, finding that this removed 61.7% of web pages. However, when filtering books (from Project Gutenberg [^5] https://www.gutenberg.org/) they first split each book by paragraph and compute the language score for each paragraph, only removing books that have an average probability of English (assigned by fastText) of below 0.5. Penedo et al. (2023) also use fastText when filtering RefinedWeb, but differently to Dolma, they use a higher threshold (0.65) and therefore filter out higher quantities of web pages.
The lower thresholds used by more recent methods are likely due to a combination of factors. First, ever-growing model sizes have necessitated larger pretraining datasets, and using a lower threshold will allow more data to pass through the filter. Additionally, it is difficult to measure the differences in accuracy across different language identification models (e.g., langdetect vs. cld3. vs. fastText) due to the differences in supported languages and strengths on text of varying lengths, but some experiments show that fastText has improved accuracy over other language identification models, [^6] https://modelpredict.com/language-identification-survey as well as reduced latency. In general, the threshold used for keeping or filtering out a data point can be in part determined by the desired dataset size, where a lower threshold will retain higher quantities of data, but possibly containing non-English text.
@@
The utility functions that have been used to detect code languages are relatively simple compared to those used for natural languages. For example, Chowdhery et al. (2023) filter for data from 24 programming languages simply by searching for documents that match a set of approved filename extension (e.g., “.py” for python). This strategy works well when filtering over data from known code data that contains file names, such as snapshots of Github repositories (Chen et al., 2021; Li et al., 2022a). However, no filtering methods that we know of find code within natural language documents, which is entirely possible within some domains such as in Stack Exchange [^7] https://archive.org/details/stackexchange and may contribute meaningful quantities of interleaved natural language and code.
@@
There are trade-offs occurring in the filter sensitivity for language detection, where a lower threshold allows greater quantities of data to pass through the filter at the risk of including lower quality data. On the one hand, using a strict filter for English data may reduce the quantities of non-English data, but on the other hand, using a lower threshold can mitigate some of the inherent biases of the language detector against dialects spoken by minority groups (e.g., African-American Vernacular English (Blodgett et al., 2016)). Determining the exact threshold to use for language classifiers is highly dependent on the use case. While fastText is the current standard for language detection due to its combination of speed and accuracy, it is not perfect. More accurate systems built in on recurrent- (Toftrup et al., 2021) and transformer-based [^8] https://huggingface.co/papluca/xlm-roberta-base-language-detection architectures have been developed, but run with higher computational costs.
Language filtering is an important first step of data selection where documents are selected to include only the desired languages. Classifier-based methods are the norm when creating both english-only and multilingual datasets. For lower-resourced languages, URL-based methods are also useful for classifying the language used in multilingual datasets and can sometimes can be more reliable than classifiers. Filtering for code languages is very simple with most works simply using the file extension. [Uncaptioned image] One interesting direction for future research is developing methods that find code data within documents of natural language. To the best of our knowledge, no such methods currently exist, but a code-language identifier could be used in domains with interleaved code and natural language to find meaningful quantities of new code data. Language Filtering Summary & Ideas
@@ 3.2 Heuristic Approaches
When pretraining a language model, raw datasets are generally composed of massive quantities of text from a wide variety of sources that are filtered through a set of simple heuristics. Major models are often trained on web scrapes such as CommonCrawl and GitHub, though transparency into their precise compositions are on the decline (Bommasani et al., 2023). It is widely known however that raw text data from the internet can contain significant quantities of boiler-plate text, error messages, and offensive text (Raffel et al., 2020; Touvron et al., 2023a; Elazar et al., 2024). For example, Gao et al. (2020) find that, when creating the Pile, the most common 13-grams were character repetitions such as a string of dashes (“– –”) with 11 million instances. Removing such undesirable text is very important, but must be done efficiently due to the size of the corpora involved. In this case, a common approach to filtering data involves simple and efficient-to-compute heuristics.
The goal of heuristic approaches is to constrain the training distribution along some dimension (e.g., sentence length, repetitiveness), with the assumption that the evaluation distribution will exhibit similar characteristics. The number of heuristics that have been used in past works is extensive, but generally fall into one of the following categories of heuristics: item count, repetition count, existence, ratio, or statistics. In this section we discuss a representative sample of them, provide Table 2 as an overview of heuristic methods, and include a comprehensive listing in Appendix A.
@@
For heuristic-based text filtering, the utility functions should be very fast to compute and generally rely on one or more qualities of the raw text included in each data point of the dataset. However, the exact text qualities that lead to best performance depend on the desired use case for the model, and are yet unknown. Thus, a wide variety of distinct utility functions have been proposed, each making assumptions about the expected testing distribution, and filtering out data that does not fit those assumptions. For example, Raffel et al. (2020) discard any web page with fewer than five sentences, Rae et al. (2021) discard documents with fewer than 50 or greater than 100,000 words, and Xue et al. (2021) discard documents with fewer than three lines of text with 200 or more characters. Another common heuristic is to remove documents with specific black-listed words or phrases. Raffel et al. (2020) remove documents with any word on the “List of Dirty, Naughty, Obscene, or Otherwise Bad Words”. [^9] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words However, Penedo et al. (2023) suggest that the NSFW word block lists as used by Raffel et al. (2020) can often lead to false positives, resulting in the over-filtering of legal and medical content.
@@
Another dimension to consider when designing a filter is the domain of interest. For example, when collecting code data for the ROOTS dataset, Laurençon et al. (2022) filter source files between 100 and 200,000 characters, with between 15-65% alphabetic characters (a-zA-Z), and line lengths between 20-1000 characters. If these filters were applied on natural language data, significant quantities of desirable data would be removed (e.g., many books have greater than 65% alphabetic characters). Additionally, Laurençon et al. (2022) filter out files that have line lengths that are highly uniform (line length standard deviation 3 characters or less). While filtering out data with uniformly lengthed lines may improve data quality for code, it would likely filter out significant quantities of poems and other rhythmic text due to the similarity in line length, and would likely remove data that has been collected from pdf files, which could be desirable when training a model for a non-code domain.
@@
Many heuristic utility functions can be used as part of a deterministic or stochastic selection mechanism. In a deterministic selection mechanism, any data points which are above/below the utility threshold will be removed. The vast majority of existing heuristic filters use a deterministic selection mechanism. However, this may unnecessarily remove data that resembles the evaluation distribution (Brown et al., 2020a). In particular, the target dataset distribution may be long-tailed in the characteristic of interest, and deterministically removing all data above/below a threshold can harm the modelling of that characteristic.
@@
One disadvantage to heuristic selection approaches is that they do not judge the quality or content of a document, rather, they rely entirely on surface level statistics and counts, lacking the “finesse” of other data selection methods. This can lead to desirable data being removed, but this is often accepted as a trade-off as long as the quantity of data after filtering is still within the desirable range.
@@
Heuristics are a very efficient method for removing large quantities of data, but lack finesse. Most common heuristic utility functions fall into 5 categories: item count, repetition count, existence, ratio, and statistics. Heuristics should be designed specifically for the domain of interest (e.g., heuristics for code may not work for natural language). Heuristics can also be used to improve data point characteristics (e.g., removing HTML from a web document, or specific lines within a document). [Uncaptioned image] Because heuristics lack finesse, it may be worth exploring the use of stochastic selection mechanisms to allow higher quantities of data through the filter (with the hope that a filter farther down the pipeline will remove data that is actually undesirable). [Uncaptioned image] Validating the quality of a heuristic filter is a very slow process, requiring either manual inspection, or training and evaluating a model. Developing methods that can directly measure the data itself could significantly improve iteration time. Heuristic Approaches Summary & Ideas
@@ 3.3 Data Quality
Training on the highest quality data can lead to stronger performance (Du et al., 2022). However, “high-quality” is not a well-defined term with respect to language model pretraining data. Additionally, there is no one-size-fits-all characterization as, depending on the use case, “high-quality” data can vary drastically (Longpre et al., 2023c). Additionally, implementations of quality filters (e.g. for detecting toxicity) can further marginalize communities whose text is not commonly considered “high quality” (Xu et al., 2021; Longpre et al., 2023c). In this work, we narrow the use of the phrase “high-quality” to a common use of the phrase referring to data that is known to be written by humans and has likely gone through an editing process (Grave et al., 2018; Gao et al., 2020; Brown et al., 2020a; Chowdhery et al., 2023). Some data domains that fall under the “high-quality” category are Wikipedia, books, patents, and peer-reviewed journal articles.
@@
One commonly used method when filtering for quality is classifier-based quality filtering, where the goal is to identify data points that are likely from the same (or similar) distribution as a known “high-quality” corpus of data points (reference corpus) (Longpre et al., 2023c). For example, Brown et al. (2020a) filters a version of Common Crawl [^10] https://commoncrawl.org/ based on similarity to “high-quality” reference corpora (including WebText (Radford et al., 2019a), Books1, Books2, and English wikipedia). They train a classifier using the “high-quality” reference corpora as the positive class and unfiltered Common Crawl documents as the negative class. This classifier then defines the utility metric for selecting data that follows the desired distribution. The classifier used to assign quality scores to each document should be fairly lightweight due to the scale of pretraining datasets. For example, Du et al. (2022) use a feature hash based linear classifier, while Brown et al. (2020a) and Gao et al. (2020) use a fastText model (Joulin et al., 2016) with an n-gram size of 2. Similarly, Xie et al. (2023b) computes the utility as an importance weight between two hashed n-gram generative models. While modern language models may be able to better model the reference corpus distribution, they would also significantly increase the amount of compute required over hashing and n-gram based methods.
@@
One potential negative impact of classifier-based quality filtering is that the reference corpora are unlikely to contain examples of all data that are considered high quality. Thus, it is desirable to allow some stochasticity in the selection mechanism. For example, Brown et al. (2020a) use their classifier to score Common Crawl documents and keep each document only if they satisfy
@@
where $\alpha=0.9$ was determined by matching the distribution of scores from the WebText corpus (Radford et al., 2019b). By selecting documents that don’t match the “high quality” distribution with some stochasticity, they ensure that documents in the final corpus are mostly high-scoring, but still include some documents that are out of distribution with respect to the reference corpora.
@@
An important consideration when performing quality filtering is that certain reference corpora may be biased towards or away from certain demographics, dialects, and socialects (Rae et al., 2021). Additionally, Gururangan et al. (2022) find that a quality filter trained on Wikipedia, books, and newswire (a replication of the quality filter from Brown et al. (2020a)), evaluated on high school newspaper articles, prefers articles from high schools in wealthier, higher-educated, and urban areas. They demonstrate that even though quality filtering is often assumed to be fairly neutral, it carries with it an implied judgment of preferred values. Thus, quality filtering should be done with care to ensure that biases are avoided and the distribution of data sufficiently covers the desired dialects and demographics. Filtering text for quality, while avoiding the introduction of biases is an important direction for future research.
@@
Data quality methods aim to select data points that are similar to data which is of known “high-quality”. Classifier- and perplexity-based quality filtering are the existing approaches, with classifier-based filtering being the more popular method. For both classifier- and perplexity-based methods, it is important to be careful when training the classifier because this can introduce biases (e.g., against lower-income, less-educated populations). [Uncaptioned image] Exactly which situations quality filtering is beneficial is still undecided, as recent works have trained performant models without using quality filters at all. This is an area that requires further study. Data Quality Summary & Ideas